Circular No. 6775

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997dt IN NGC 7448 Q. Y. Qiao, Y. L. Qiu, W. D. Li, and J. Y. Hu, Beijing Astronomical

Observatory (BAO); and A. Esamdin, Urumqi Astronomical Station, report the discovery of a supernova (mag ∼ 15.3) on unfiltered CCD images taken with 0.6-m telescope on Nov. 22.44 UT. The supernova is located at α = 23^{h}00^{m}02^{s}.93, δ = +15^{o}58′50′′.9 (equinox 2000.0), which is 9′′.3 west and 1′′.0 north of the center of NGC 7448. SN 1997dt was confirmed on images taken on Nov. 21.48 (mag ∼ 15.4). Unfiltered CCD images of the same field taken on Nov. 14 show no star at the position of the supernova (limiting mag ∼ 18.5). A low-dispersion spectrum of the supernova, obtained on Nov. 22.50 by A. Esamdin, J. Y. Wei, L. Cao (BAO), and Q. S. Gu (Nanjing University) with the 2.16-m telescope at Xinglong Station, shows this to be a type-Ia supernova near maximum. The spectrum displays a strong Si II 615-nm absorption feature. The recession velocity, derived from the narrow Hα emission line, is 2192 km/s.

GS 1354–64 A. J. Castro-Tirado, Laboratorio de Astrofísica Espacial y Física Fun-

damental, Madrid; S. Ilovaisky, Observatoire de Haute-Provence; and H. Pedersen, Copenhagen University Observatory, write: “Following the detection of GS 1354–64 by Rossi XTE (IAUC 6772) and BATSE (IAUC 6774), B- and R-band images were obtained with the New Technology Telescope at the European Southern Observatory on Nov. 22.35 UT by J.-F. Gonzalez, M. Pizarro, J. Miranda, and H. Boehnhardt. The optical counterpart (cf. Kitamoto et al. 1990, Ap.J. 361, 590) is being detected in a bright state. Preliminary magnitudes: R = 16.9 ±0.1, B = 18.1 ± 0.2.”

PSR J0633+1746 Yu. P. Shitov and V. M. Malofeev, Pushchino Radio Astronomy Ob-

servatory, Lebedev Physical Institute, communicate: “With our colleagues O. I. Malov and V. D. Pugachev, we observed Geminga at low frequencies with the east-west arm of Wide-Band Cross-type Radio Telescope DKR-1000 (area ∼ 40 000 m^{2}) during July–October and detected radio emission of the pulsar PSR J0633+1746 with period 0.237 s at 41 and 61 MHz (cf. IAUC 6559). The flux density is ∼ 300 mJy at 41 MHz, and the spectral index > 2.5, suggesting a mean flux density < 30 mJy at 102.5 MHz. The width of the integrated profile at 50 percent of maximum intensity is about half the period at both frequencies.”

1997 November 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green